OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f of x as x approaches 5 where f of x equals 5 minus x when x is less than 5, 8 when x equals 5, and x plus 3 when x is greater than 5 0 8 3 The limit does not exist.

OpenStudy (anonymous):

@Nerdgirl

OpenStudy (anonymous):

can you help me with this one too?

OpenStudy (nerdgirl):

Hit me.

OpenStudy (nerdgirl):

I'll certainly do my best.

OpenStudy (anonymous):

thank you

OpenStudy (nerdgirl):

You're welcome. What's the question?

OpenStudy (anonymous):

i need to find the limit for the equation

OpenStudy (anonymous):

lim f(x), f(x)= x-5 x<5 8 x=5 x+3 x>5

OpenStudy (anonymous):

thats the equation

OpenStudy (anonymous):

my teacher gave us notes on this, but i didn't understand the lesson very well

OpenStudy (zarkon):

\[f(x)=\left\{\begin{array}{ll} 5-x & ,x<5 \\ 8 &, x=5 \\x+3 &,x>5 \end{array}\right.\]

OpenStudy (anonymous):

yes thank you @Zarkon

OpenStudy (jdoe0001):

hmm right \(\large { lim_{x\to 5}\ f(x) \begin{cases} 5-x&x<5\\ 8&x=5\\ x+3&x>5 \end{cases} }\)

OpenStudy (anonymous):

yup

OpenStudy (jdoe0001):

well... the issue is a matter of continuity notice the 1st equation, the one on the "left side" of 5, what's its limit? notice the 2nd, is just 8, so is a fixed value notice the 3rd one, what's its limit we know that when x = 5, f(x) = 8 so for the continuity to be maintained, the other 2 fellows, to the left and right side of 5 should give 8 also, so the line is continuous and if it's so, then the "double-sided limit" does exist

OpenStudy (anonymous):

so it doesn't exceed the limit?

OpenStudy (nerdgirl):

Correct.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

but yes... the x+ 3 will give a limit of (5)+3 = 8 the 5-x will give a limit of 5-(5) = 0 those 2 fellows, the right-hand-side equation and the left-hand-side equation do not meet at 8, thus the double-sided limit doesn't exist

OpenStudy (anonymous):

ok thank you

OpenStudy (jdoe0001):

yw